Solve the system using substitution or elimination:

3x+2y= -12
2x+y= -7
and
2x-4y= -2
4x+8y=0

Answer by checkley71(8403)

3X+2Y=-12 NOW MULTIPLY THE OTHER EQUATIOBN BY -2 & ADD THEM
-4X-2Y=14
--------------------------
-X=2
X=-2 ANSWER.
3*-2+2Y=-12
-6+2Y=-12
2Y=-12+6
-2Y=-6
Y=-6/2
Y=-3 ANSWER.
PROOF
2*-2-3=-7
-4-3=-7
-7=-7
-------------------------------------------
MULTIPLY THE FIRST EQUATION BY 2 & ADD THEM
4X-8Y=-4
4X+8Y=0
----------------
8X=-4
X=-4/8
X=-1/2 ANSWER.
4*1/2+8Y=0
2+8Y=0
8Y=-2
Y=-2/8
Y=-1/4 ANSWER.
PROOF
2*-1/2-4*-1/4=-2
-1-1=-2
-2=-2